OpenStudy (nuccioreggie):

Multiply two and four-fifths negative five and two-thirds . negative ten and eight-fifteenths negative fifteen and thirteen-fifteenths ten and eight-fifteenths fifteen and thirteen-fifteenths

OpenStudy (some.random.cool.kid):

easy just multiply it ;]

OpenStudy (nuccioreggie):

i thought it would be b

OpenStudy (nuccioreggie):

@some.random.cool.kid

OpenStudy (some.random.cool.kid):

yeah :} its not c or d so yeah

OpenStudy (some.random.cool.kid):

gtg

OpenStudy (nuccioreggie):

@DanJS

OpenStudy (nuccioreggie):

@TuringTest

OpenStudy (danjs):

hi

OpenStudy (nuccioreggie):

hi can u please help me with this i thought it was b @DanJS

OpenStudy (danjs):

\[2\frac{ 4 }{ 5 } * -5\frac{ 2 }{ 3 }\]

OpenStudy (nuccioreggie):

yea

OpenStudy (danjs):

First i would change each mixed number into improper fractions

OpenStudy (danjs):

sorry , the site is lagging like crazy

OpenStudy (nuccioreggie):

its ok

OpenStudy (danjs):

\[\frac{ 5*2 +4 }{ 5 } * \frac{ -5*3+2 }{ 3 }\]

OpenStudy (danjs):

you know how to turn them into improper fractions like that?

OpenStudy (nuccioreggie):

no but is the answer c dan?

OpenStudy (nuccioreggie):

@DanJS

OpenStudy (danjs):

you dont care how to do it, just want the answer?

OpenStudy (nuccioreggie):

yea i care but im trying to jus hurry n get it done so i can go see my aunt in the hospital thats all